Shift Summary - Evening Jan 19, 2000


Subject: Shift Summary - Evening Jan 19, 2000
From: Vivian O'Dell (odell@fnal.gov)
Date: Fri Feb 04 2000 - 15:33:07 CST


              Shift Summary for Evening shift, Wed. 1/19/2000

On Shift: RR, DJ and RK + ...

Hours of beam available at all intensities: 5.0/8.5
Hours of beam available at nominal* intensity: 5.0
Hours of KAMI test data: 4.5
Hours of special runs (laser scans, muon runs...): 0.0

*For these purposes, nominal is whatever we rewqusted.
 
Accounting of lost time & intensity:

  One quench
  Intensity Changes
  starting/stopping runs
  Two accesses to NM2, .5 hr cool-off for each.
 
========================================================================
Runs ended this shift:

       15620 to 15629 -- test KAMI_HADRON
       Tapes: QKE210-218, QKB170 for runs 15620-15629

========================================================================
Tapes brought to FCC this shift: None.
========================================================================
Accesses during shift:
     NM2 to reinstall the shadow absorber
     NM2 to install 44.5 inches of polyethylene betweeen NM2S3 and NM2SR
========================================================================
Overview of shift:

       o KAMI test running in HADRON mode.
         Measurement of the neutron/kaon ratio
         (using various absorber thickness) with the UCAL.
       o Measure the kaon flux using KL->3pi0

========================================================================
Description of Shift

 15:25 Had called to start cool-down for NM2 access
 15:30 'official shift change'
 16:16 TeV Quench
 16:45 RC and DJ return from ABS installation
 17:20 15620 Get 5E11, 38.5" of Be in
 18:42 15621 2E11 38.5" Be
 18:51 15622 2E11 38,5"

 19:20 Plan access to NM2 to install polyethylene
 20:15 15625 2E11, 44.5" polyethylese, all Be
               The polylethylene is in the form of 5 hunks on a U-channel
               tray between NM3S3 (aka musweep2 ) and NM2SR (aka spin rotator)
 20:54 16626 5E11, poly 38.6"" of Be
 22:15 16626 end the run, Remove NM2ABS1, > 20.6" Be + Poly
 22:25 16627 Start Run
 23:10 Stop run Go to NM2ABS2 to position 2 => 10.3" Be
 23:20 16628 Start Run
 23:40 Stop run. NM2ABS2 => "0", => No Be Absorber
 23:45 16629 Start new run. No Be.
 23:59+ Finish the last of these runs....
 
 Details of the runs may be found in the file:
      da_ktev/doc/runkami.log
  ............................................................................
  Summary of on-line UCAL/Det(3pi0) ratios (using ucal#nkratio) [on-line]:
  (continue with table from previous shift)
       
            Absorber UCAL(E>30GeV)/Det(3pi0)
          ----------------------------
               0" Be 1.51
              10" Be 1.13
              20" Be 1.00 (by defintion)
              38" Be .78
         poly+38" Be .62
         poly+20" Be .75
         poly+10" Be .85
         poly+ 0" Be .96

========================================================================
Follow up Notes on Previous Shift:
   Collected more AGEL data on floppy disks - until about 8:30 pm.
========================================================================
Items to Follow up on Next Shifts:
   Thursday AM starting around 7-8 am, Rick Ford et al will do a SEM
   calibratoin
========================================================================

  Douglas Jensen DJensen@fnal.gov x8194



This archive was generated by hypermail 2b27 : Fri Feb 04 2000 - 15:33:20 CST